how to solve this: f(x)= x2+x and g(x)=x2-x, question (f.g)(x) =?
Answer by stanbon(75887)

how to solve this: f(x)= x^2+x and g(x)=x^2-x, question (f.g)(x) =?
--------------------------------
fog(x) = f[g(x)] = f[x^2-x] = (x^2-x)^2 + (x^2-x)
= x^4 - 2x^3 + x^2 + x^2 - x
---
= x^4 -2x^3 + 2x^2 - x
===========================
Cheers,
Stan H.